Filer Manufacturing has 6,726,499 shares of common stock outstanding. The current share price is $31.1, and the book value per share is $5.8. Filer Manufacturing also has two bond issues outstanding. The first bond issue has a face value of $51,715,711, has a 0.05 coupon, matures in 10 years and sells for 83 percent of par. The second issue has a face value of $64,694,774, has a 0.06 coupon, matures in 20 years, and sells for 92 percent of par.

The most recent dividend was $2.91 and the dividend growth rate is 0.07. Assume that the overall cost of debt is the weighted average of that implied by the two outstanding debt issues. Both bonds make semiannual payments. The tax rate is 0.27.

What is Filer's aftertax cost of debt? Enter the answer with 4 decimals (e.g. 0.2345)
Answer #1

First Bond Issue(B1):

To find the YTM, we need to put the following values in the financial calculator:

N = 10*2 = 20;

PV = -0.83 * 51,715,711 = -42,924,040.13;

PMT = (0.05/2)*51,715,711 = 1,292,892.78;

FV = 51,715,711;

Press CPT, then I/Y, which gives us 3.72

Periodic Rate = 3.72%

So, YTM = Periodic Rate * No. of compounding periods in a year = 3.72% * 2 = 7.44%

Second Bond Issue(B2):

To find the YTM, we need to put the following values in the financial calculator:

N = 20*2 = 40;

PV = -0.92 * 64,694,774 = -59,519,192.08;

PMT = (0.06/2)*64,694,774 = 1,940,843.22;

FV = 64,694,774;

Press CPT, then I/Y, which gives us 3.57

Periodic Rate = 3.57%

So, YTM = Periodic Rate * No. of compounding periods in a year = 3.57% * 2 = 7.13%

Weight of B1 = B1 / (B1 + B2)

= 42,924,040.13 / (42,924,040.13 + 59,519,192.08) = 0.4190, or 41.90%

Weight of B2 = B2 / (B1 + B2)

= 59,519,192.08 / (42,924,040.13 + 59,519,192.08) = 0.5810, or 58.10%

Weighted Average Cost of Debt = [wB1 * kB1] * [wB2 * kB2]

= [0.4190 * 7.44%] + [0.5810 * 7.13%] = 3.12% + 4.14% = 7.26%

After-tax cost of debt = Weighted average cost of debt * (1 - t)

= 7.26% * (1 - 0.27) = 5.30%, or, 0.0530